OpenStudy (anonymous):

Please help me check my work: Solve for w A=2L+2w A-2L=2L+2w-2L A-2L=2w ---Divide each side over 2 A-L=w Where is my mistake?

OpenStudy (anonymous):

A=2L+2W 2L+2W=A -2L -2L 2W= A-2L divided by 2 on both sides w= (A-2L)/2

OpenStudy (anonymous):

You can also right it as w= (1/2)(A-2L)

OpenStudy (anonymous):

Makes sense now, of course, lol. Thanks!

OpenStudy (anonymous):

I always like to switch the side around which I am solving for to the left so I follow PEMDAS